OpenStudy (anonymous):

Let U denote the universe and A, B, and C be three subsets. Prove that if A u B = A^c u C, then B u C = U. (I'll rewrite it in comments)

OpenStudy (anonymous):

If \[A \cup B = A^c \cup C,\] then \[B \cup C = U\]

OpenStudy (anonymous):

i think you're missing a piece of info...

OpenStudy (anonymous):

are A, B & C disjoint?

OpenStudy (anonymous):

All it told me was that...that would help though, wouldn't it? I think just assume theyre not.. Does that make it impossible though?

OpenStudy (zarkon):

you have all the information you need to do the problem

OpenStudy (anonymous):

So do you know how to do it?

OpenStudy (zarkon):

I'll get you started let \(x\in U\) then \(x\in A\) or \(x\in A^c\) assume \(x\in A\) then \(x\in A\cup B\) therefore \(x\in A^c\cup C\) since \(x\in A\) we know that \(x\notin A^c\), therefore \(x\in C\) now show what happens if we assume \(x\in A^{c}\)

OpenStudy (anonymous):

How do you know \[x \in A^c \cup C\] from \[x \in A \cup B\]?

OpenStudy (zarkon):

you are given \[A \cup B = A^c \cup C\]

OpenStudy (anonymous):

Oh, duh...thanks lol

OpenStudy (anonymous):

So now assume \[x \in A^c\] therefore \[x \in B\] since \[x \notin A^c.\] So \[x \in A \cup B\] I dont know if I'm on the right track...

OpenStudy (anonymous):

*since \[x \notin A\]

OpenStudy (zarkon):

Assume \(x\in A^c\) then \(x\in A^c\cup C\)...

OpenStudy (anonymous):

Okay thank you. My teacher sucks lol